OpenStudy (anonymous):

Can anyone please explain how to solve for y and x through substitution and elimination? If possible, step by step please. You can use this as your example: 2y - 3x = 30 4y + 3x = 24

OpenStudy (anonymous):

Add the equations 6y = 54 y = 9

OpenStudy (mimi_x3):

elimination: 2y - 3x = 30 - 4y+ 3x = 24 -2y = 6 y = -3 To find x sub it into 2y - 3x = 30

OpenStudy (anonymous):

:< I'm afraid I don't quite understand, Ishaan.

OpenStudy (mimi_x3):

lols , btw , yu cant add the 2 equations in elimination xDD

OpenStudy (anonymous):

2y - 3x = 30 4y + 3x = 24 Now add these equations 2y + 4y +3x -3x = 30 + 24 6y = 54 y = 9

OpenStudy (anonymous):

Thanks, Mimi. (: I understood it! ^(OwO)^ But how about through substitution? o3o

OpenStudy (mimi_x3):

ohh yeah my bad

OpenStudy (mimi_x3):

i read the equation wrongly xDD

OpenStudy (anonymous):

2y -3x = 30 2y - 30 =3x This is the fist equation now second 4y + 3x =24 Now substitute the value of 3x 4y + 2y -30 = 24 6y =54 y=9

OpenStudy (mimi_x3):

2y - 3x = 30 + 4y + 3x = 24 6y = 54 y = 9

OpenStudy (anonymous):

Now substitute y in first 2x9 -3x = 30 18-30 = 3x 6-10=x x=-4

OpenStudy (mimi_x3):

lols , my way is the elimination method btw

OpenStudy (anonymous):

how do u know when to use substitution and when to use elimination

OpenStudy (anonymous):

spicenangel depends on the equation. if i see same co-efficient then I eliminate or if see some multiplication that makes co-efficient same. elimination is faster method you will get to know after enough practice.

OpenStudy (anonymous):

So, it's like this, through elimination, based on the site you gave me, and info you guys gave me, and from what I know? eliminate x 2y - 3x = 30 +4y + 3x = 24 6y + 0x = 54 6y + 0x = 54 6y = 54 Then, you divide those two by 6, to erase the 6 from the y. y = 9 eliminate y 2y - 3x = 30 Replace y with your y with the y at the first equation. 2(9) - 3x = 30 18 - 3x = 30 Then you transpose 18 to the other side because you need to have one side with all the numbers with variables, and the other side with numbers only. ((18 becomes -18 because it was transferred to the other side, right?)) -3x = 30 - 18 And... this was where I was stumped. Do you divide the two with -3 already or do you subtract 30 and 18 first?

OpenStudy (mimi_x3):

-3 x = 12 -x = 4 x= -4

OpenStudy (anonymous):

Ah, so I subtract the two first. Thanks. owo